OpenStudy (anonymous):

Ln(X+1)^2=2 Please solve and explain how. I will medal.

OpenStudy (kinggeorge):

Is the Ln the natural log?

OpenStudy (anonymous):

properties of logs: so, 2ln(x+1) = 2 ln(x+1) = 1 e = x + 1 x = e-1

OpenStudy (anonymous):

^ Exactly what Gralwyn said :P

OpenStudy (kinggeorge):

@Gralwyn The way I'm interpreting it, \(\ln(x+1)^2\neq2\ln(x+1)\). The log is squared, and not the \(x+1\).

OpenStudy (anonymous):

@KingGeorge Then it would be written as (ln(x+1))^2

OpenStudy (anonymous):

\[(\ln(x+1))^2\]

OpenStudy (kinggeorge):

I can't distinguish between either in this case. Do you know which one it is @Bigjoe11?

OpenStudy (anonymous):

But the way it is written, it doesn't have those parantheses, so it can't be the whole log squared.

OpenStudy (anonymous):

Even if it wasn't, properties of logs would still allow the 2 to be moved to the front. And like Gralwyn said, the whole quantity including the log would have to be squared

OpenStudy (anonymous):

@dixiemitsy are you sure about that?

OpenStudy (kinggeorge):

The properties of logs would not allow the 2 to be moved in front. If the whole log were squared, this would actually become a rather difficult question. In retrospect, you are probably correct Gralwyn. But ordinarily, I would interpret the logarithm being squared.

OpenStudy (kinggeorge):

In fact, if you type the problem into wolfram as is, wolfram will interpret it as the logarithm being squared.

OpenStudy (anonymous):

Lets say that the whole log were squared, then the answer would be: ln(x+1) = sqrt(2) e^(sqrt(2)) = x+1 e^(sqrt(2)) - 1 = x e^(2^(1/2)) - 1 = x e - 1 = x what?

OpenStudy (anonymous):

i probably did something dumb when i multiplied the 2 and 1/2 since e^(sqrt(2)) isn't equal to e^1

OpenStudy (anonymous):

Still never understood why I can't do that

OpenStudy (kinggeorge):

The 2 and power of 1/2 don't quite cancel like that. But other than that it looks mostly fine. You would get a second solution of \(x=e^{-\sqrt2}-1\) as well though.

OpenStudy (kinggeorge):

So depending on what the exact question is, the solution should either be\[x=e-1\]or\[x=e^{\sqrt2}-1,\quad x=e^{-\sqrt2}-1\]

OpenStudy (anonymous):

actually when you write ln(x+1)^2, this actually means [ln(x+1)]^2. If you want (x+1) to be squared, you must be specific and write ln[(x+1)^2]

OpenStudy (anonymous):

the same convention holds for sin(x)^2 = sin^2(x) = (sinx)^2

OpenStudy (anonymous):

Of course, not all people follow this convention. This why it is always better to use parentheses. Plus calculators are also programmed to follow this convention.

OpenStudy (kinggeorge):

Exactly my thoughts. Well said.

OpenStudy (anonymous):

so if you type ln(2)^2, your calculator will compute (ln(2))^2

OpenStudy (anonymous):

Thanks for the help guys! It helped a ton!